# The Superclouds of the Local Milky Way

### v1

@@ 3.2.2 Shape Definition
To describe the 3D shape, we apply principal component analysis (PCA; Jolliffe 1986 ) to each individual structure. PCA identifies the direction of largest variance within a dataset, determining the primary orientation of the structure. In the case of 3D data, PCA extracts three mutually orthogonal components, creating a new set of basis vectors that aligns with the intrinsic shape of the structure. The first principal component (PC1) describes the direction of the largest variance and the length of the structure. The second principal component (PC2) describes the direction of the largest variance orthogonal to PC1 and the width of the structure. The third principal component (PC3) describes the direction of the largest variance orthogonal to PC1 and PC2, and the height of the structure.
We determine the structure axis lengths by projecting the data onto the principal components and measuring the extent between the minimum and maximum, resulting in an upper boundary. For the length, this is a good approximation, but due to the irregular shape of the clouds, defining the height and width leads to a large overestimation in the majority of the cases.
@@ 3.2.3 Orientation Calculation
The structure orientation is determined by the angle between PC1 and a chosen reference axis. This angle is calculated by applying the arccosine to the absolute value of the cosine similarity between the two vectors, and then converting the result to degrees. For consistency and easier comparison, all structures are defined as orientated in the positive direction relative to the reference axis, with angles measured clockwise from it. The angle between PC1 and the Y-axis (face-on) is defined as $\alpha$, while $\beta$ represents the angle between PC1 and the Z-axis (edge-on). Circular statistics are applied to calculate the mean and standard deviation of all angles $\alpha$ and $\beta$. The angles are weighted by mass to consider each structure’s relative importance.
@@ 3.2.4 Undulation Fitting
where $z$ corresponds to the projected Y-coordinate, obtained by rotating the superclouds anticlockwise by their average angle $\alpha$ to lie in the plane, and $f(z)$ gives the corresponding Z-position. The parameter $a$ is the amplitude, $\omega$ the angular frequency, $\phi$ the phase shift, and $d$ the vertical offset. The function includes an exponential modulation of both amplitude and frequency: $\epsilon_{a}$ describes the exponential growth or decay of the amplitude along the wave, while $\epsilon_{\omega}$ describes a linear change in frequency. Positive values for $\epsilon_{a}$ indicate a growth in amplitude over distance, while negative values indicate damping. For $\epsilon_{\omega}$, a positive value indicates a growth in frequency, resulting in a decrease in wavelength via the relation $\lambda=2\pi/\omega$, where $\lambda$ is the wavelength and $\omega$ the angular frequency.
The fitting approach is simplified to give an initial characterization of the underlying undulation, and should not be interpreted as implying a physical oscillation.
@@ 4.1 HOP Cloud Catalog
HOP identifies 40 clouds in the E24 dust map, visualized in the left panel of Fig. 1. The cloud masses range from $3.8\times 10^{4}$ to $2.8\times 10^{6}$ M, a selection of the calculated cloud properties including the mass and orientation can be found in Table 2.
Figure 2: Orientation distribution of the 40 HOP clouds, the marker color corresponds to the same cloud as in Fig. 1. Distance from the center reflects the cloud mass, the solid line represents the mean orientation. Top: face-on view from the North Galactic Pole, with an average angle of $\alpha=22.6\pm 25.5\degr$. Bottom: edge-on view from the Galactic Center, with an average angle of $\beta=92.3\pm 23.4\degr$.
@@
Figure 3: Face-on view of the HOP clouds color-coded by their mean altitude in Z. The Sun is represented by the yellow dot in the center, the Galactic Center is located to the right.
@@ 4.2 Supercloud Catalog
Taking the orientation, proximity, and 3D alignment of the HOP clouds into account, we identify seven superclouds, visualized in the left panel of Fig. 4. The corresponding HOP cloud components are listed in Table 1.
To connect the superclouds, the first criterion taken into account is the face-on orientation of the HOP clouds, defined by $\alpha$ and shown in the right panel of Fig. 1. If adjacent clouds show a similar orientation, they are considered as initial candidates for a supercloud. This is verified considering their proximity, candidate clouds that are spatially connected are confirmed as parts of the superclouds. Clouds that deviate from the group orientation, yet still show a spatial connection, are also included as candidate clouds. It is likely that external forces have played a role in the fragmentation of the superclouds during their lifetime. Clouds showing no connection to the supercloud candidates are verified concerning their general 3D alignment. If displaced clouds follow the spatial trend of a larger, already verified group, they are accepted candidates based on their 3D alignment. If these clouds cannot be assigned to another candidate group following the mentioned steps, they are confirmed as part of the first suggested supercloud.
Figure 4: Face-on view of the seven superclouds defined in the local Milky Way (left) and their corresponding orientation (right). Each cloud is assigned a distinct color, the length of the line corresponds to the calculated length of the cloud. The Sun is represented by the yellow dot in the center, the Galactic Center is located to the right. An interactive 3D version of the seven superclouds is available here.
@@ 4.2.1 Supercloud Identification
For the more disrupted orange and yellow cloud group, we adopt the name Malpolon cloud. Its irregular morphology and puffed-up appearance in Z is likely caused by the influence of nearby high-mass stars, as traced by Zari et al. (2023) and the ALS III Catalog of Galatic OB stars; Pantaleoni González et al., under review, which is a recent expansion and Gaia DR3 update of the catalog in Pantaleoni González et al. (2021).
The green supercloud, which we name the Vela Ridge cloud, is the least spatially coherent among the identified superclouds, with its components fragmented and widely separated. A possible explanation is that Vela Ridge itself is not an independent supercloud, but rather a disrupted part of the Sagittarius Spur Extensions, as orbital tracebacks of young star clusters hint that past clustered supernovae may have occurred in this region (Swiggum et al., in prep). This theory of a past connection is also supported by the presence of high-mass stars in the canyon that separates clouds 4 and 33 from cloud 2. The stellar component extends the Spur further to almost double the initially defined length, linking the separated cloud components (Pantaleoni González et al., under review). Spatially, it is not unlikely that clouds 29 and 39 (referred to as the L-clouds) are being sheared off from the SSE due to external forces, such as feedback from high-mass stars in this region (Zari et al. 2023, Pantaleoni González et al., under review), or supernovae from the Messier 6 cluster family that partially occupies this region (Swiggum et al., 2024). Additionally, clouds 23 and 26 may have originally been part of the Split, which is supported by their physical connection. The displacement is likely a result of the event responsible for the ring they surround. We keep Vela Ridge as a supercloud in this work, as we cannot verify any of the alternative interpretations yet.
Lastly, the brown supercloud, located at the upper edge of the map, is given the name Anguis cloud. While the two supercloud components are not connected, comparing our results with the extended version of the Edenhofer et al. (2024c) dust map and the dust map by Vergely et al. (2022) shows that they are likely part of a supercloud extending beyond the map.
@@ 4.2.2 Supercloud Properties
All superclouds we identify are truncated by the artificial boundary of the 3D reconstruction and are therefore incomplete. Their cloud components are either cut off at the border of the map, or show potential extensions when compared to the 2 kpc dust map by Edenhofer et al. (2024c) and the dust map by Vergely et al. (2022). These maps are not used for the segmentation itself, due to their decreasing reliability beyond the boundary of the main map and their lower resolution. Instead, they serve as additional references for assumptions on the possible continuation of the superclouds.
As shown by Alves et al. (2020), the Radcliffe wave extends to the Cygnus X complex. Zucker et al. (2020) calculated accurate distances for SFRs in the Star Formation Handbook (Reipurth, 2008a, b) up to 2.5 kpc from the Sun. Comparing their results with the superclouds, we find that the Natrix cloud is likely to extend to the Rosette Nebula, while the SSE connects to M20, Lagoon, L291, NGC6604, M17, M16, and L379. It is interesting to note that each supercloud, excluding the Anguis cloud, contains at least a few known star-forming regions. This absence could indicate a less active or earlier phase of cloud evolution, but it might also result from observational restraints or incomplete coverage of the region.
The masses of the inferred superclouds range from $1.1$ to $3.6\times 10^{6}$ M, with the Anguis cloud as an outlier at $7.8\times 10^{5}$ M. All superclouds, except the Anguis cloud, show lengths ranging from approximately 1.5 to 2.5 kpc, which is calculated using PCA and taking all cloud components into account. A selection of the calculated cloud properties including mass and orientation can be found in Table 3. Their mean orientation is characterized by a face-on pitch angle $\alpha=33.5\pm 4.0\degr$, and edge-on angle $\beta=87.7\pm 4.2\degr$, making an association with a major spiral arm unlikely (Franco et al., 2002; Efremov, 2009). We show the face-on orientation of the superclouds in the right panel of Fig. 4. The identified superclouds are consistent with the characteristic scales and masses predicted in earlier studies (Elmegreen & Elmegreen, 1987; Park et al., 2023), representing the largest scale of ISM structures in the local Milky Way and likely throughout the Galaxy as a whole.
In total, the E24 dust map has a gas mass of $3.3\times 10^{7}$ M. For Z = ± $\pm\,100$ pc, the volume filling factor of the superclouds is 9.94%, containing 55.95% of the gas mass in the Galactic plane.
@@ 4.2.3 Undulation Pattern
Notably, the undulation observed in the Radcliffe wave is not unique. All but one of the seven superclouds show an undulating pattern in the Z/Y plane, indicated by the color pattern of Fig. 3, and are likely oscillating (Konietzka et al., 2024). Here, the upper part of the Radcliffe wave appears flat in Z, as a result of averaging the change in height across the entire length of HOP cloud 0, erasing the underlying pattern.
In Fig. 5 we show the sinusoidal fit for the selected superclouds: the Malpolon cloud, Natrix cloud, Radcliffe wave, and Vela Ridge cloud. The SSE and the Anguis cloud are excluded from the analysis, as they show a vertical displacement, but it is not as prominent as for the previous examples. The Split is also excluded, as it appears unusually flat, with only two noticeable downward extensions at Y = -200 pc (referred to as the ”C” and discussed in Edenhofer et al. 2024a) and Y = 500 pc. All resulting fit parameters are listed in Table 4, the fitting approach is simplified to give an initial characterization of the underlying undulation.
For the Malpolon cloud, we exclude the three highest altitude clouds (cloud 10, 16, and 17) while fitting. It is likely that these clouds were pushed up by nearby high-mass stars (Zari et al. 2023, Pantaleoni González et al., submitted to MNRAS), and therefore distort the inherent structure of the supercloud. The Malpolon cloud, Natrix cloud, and Radcliffe wave are well described by a damped sinusoidal wave. For the Malpolon cloud, the parameters indicate decay in the amplitude and growth in the wavelength. Both damping parameters for the Natrix cloud indicate growth over distance. For the Radcliffe wave, the parameters show decay in amplitude and wavelength.
Vela Ridge is fitted with a simple sinusoidal function. While the undulation is visible, the large wavelength is likely a consequence of the fragmented nature of the supercloud. This, together with the incompleteness, support the idea that Vela Ridge may not be an individual supercloud.
Figure 5: 2D projection of selected superclouds in the Z/Y plane with a sinusoidal fit. For the fit, the clouds are rotated anticlockwise by their $\alpha$ in the X/Y plane. The first three clouds are fit with a damped sinusoidal, the fourth cloud with a simple sinus function.
@@ 5 Discussion
A key implication of the main results in this paper is that these superclouds are unlikely to be associated with the spiral arms of the Milky Way, which have significantly smaller pitch angles (on average 10, as shown in Reid et al. 2019). This includes the most massive of the local superclouds, the Radcliffe wave, normally associated with the Local Arm (also known as the Orion Arm or the Orion Spur), e.g., Swiggum et al. (2022). Instead, the local superclouds seem to be part of a distinct, coherent structure that may reflect larger-scale dynamical processes in the Milky Way, perhaps akin to the high-pitch-angle gas lanes observed in other galaxies.
@@
Additionally, given that none of the seven superclouds seems to be associated with a Milky Way spiral arm, this implies that there is no obvious external trigger for the formation of GMCs inside them, like a spiral shock. An initial comparison with the spatial distribution of known star-forming GMCs shows that nearly all lie within these superclouds, with many concentrated along their central regions of higher density, a result aligned with earlier results from Dame et al. (1986); Elmegreen & Elmegreen (1987). This result supports a molecular cloud formation scenario in the local Milky Way in which H2 forms in the denser regions of the neutral atomic cold medium without an obvious external trigger mechanism. This is tentative at this point, and more work is needed to qualify this statement. Since GMCs are the primary sites of star formation, the superclouds presented in this work are likely to play an indirect but critical role in shaping the star formation process in the local Milky Way.
@@
Of the seven superclouds identified in this study, only two were previously known: the Split (Lallement et al., 2019) and the Radcliffe wave (Alves et al., 2020). While the Split lies on the Galactic midplane, the Radcliffe wave surprisingly undulates above and below it. Recent work by Konietzka et al. (2024), using young stellar clusters, confirmed that the undulations of the Radcliffe wave represent vertical oscillations relative to the plane. Our results show that such undulations are not unique to the Radcliffe wave. In fact, all superclouds, except the Split, exhibit similar vertical displacements. The origin of these undulations remains uncertain. One possibility is that they could result from shocks or gravitational perturbations as superclouds interact with spiral arms in the past, but further investigation is needed to test this scenario.
@@
The identified superclouds can be associated with multiple known structures in the solar neighborhood. Spatially, the Anguis and the Malpolon cloud appear to be a small section of the Cepheus spur (Pantaleoni González et al., 2021). This Spur extends from Cygnus OB2 towards the Galactic anti-center, as traced by young stellar populations and in more extended dust maps (e.g., Lallement et al., 2019). The bottom part of the Radcliffe wave (cloud 27, 34, 30, 12, and 18) and of the Vela Ridge cloud (cloud 21 and 2) define the boundaries of the GSH 238+00+09 supershell (Heiles, 1998). This kiloparsec long dust shell likely formed from clustered supernovae associated with the Cr135 cluster family (Swiggum et al., 2024), and occupies the full width between the two superclouds. Recently, Soler et al. (2025) calculated the kinetic energy densities of the local ISM based on the E24 dust map. Kinetic energy overdensities represent areas that differ roughly by a factor of ten or more from the mean value. The largest of these overdensities can be connected to the Radcliffe wave and the Split, likely as a result of past supernovae or large-scale galactic dynamics.
@@
Figure 6: Properties calculated for the seven superclouds. The color corresponds to the same cloud as in Fig. 4. RadWave refers to the Radcliffe wave, SSE to the Sagittarius Spur Extension. Top: total mass of the superclouds. Second: total length of the superclouds. Third: linear mass of the superclouds, the ratio of total mass to total length. Bottom: average density of the superclouds. The trends represent a lower boundary, as all superclouds are incomplete and likely to continue beyond the map.
The well-defined, quasi-parallel pattern of superclouds shown in Fig. 4 suggests a relatively undisturbed local ISM. While stellar feedback is clearly reshaping the gas on kiloparsec scales, for example, in the case of the Heiles supershell (Swiggum et al., 2024), it does not appear to have significantly disrupted the overall supercloud pattern. Based on the formation scenario for GMCs discussed above, we can make a conservative, lower age estimation by considering the ages of the oldest stars formed within the superclouds. For the Radcliffe wave, this corresponds to approximately 20-30 Myr. Compared to kinematic timescales, this lower age estimate is short. Recent measurements of the local ISM velocity dispersion, approximately $11\,\mathrm{km\,s^{-1}}$ (see Soler et al., 2025, and references therein), suggest that a perturbation would require over $90\,\mathrm{Myr}$ to cross the full length of a supercloud. This timescale is about three times longer than the estimate based on stellar ages and exceeds the typical lifetimes of CO clouds (see, for example Miville-Deschênes et al., 2017; Chevance et al., 2020). These discrepancies imply that superclouds are unlikely to form by a bottom-up process, i.e., from small to large scales, as would be expected in a scenario driven purely by stellar feedback. Instead, large-scale Galactic dynamics and instabilities probably play a central role in the formation of the kiloparsec-scale density structures we identify as superclouds (see, for example Kim & Ostriker, 2002, 2017; Smith et al., 2020; Colman et al., 2022). This does not exclude the influence of energy and momentum injection from interior star formation (Swiggum et al., 2024), but it points to the need for additional physical processes to explain the observed coherence and scale of these structures.

### v3

@@ 3.2.2 Shape definition
To describe the 3D shape, we applied principal component analysis (PCA; jolliffee1986pca) to each individual structure. PCA identifies the direction of largest variance within a dataset, determining the primary orientation of the structure. In the case of 3D data, PCA extracts three mutually orthogonal components, creating a new set of basis vectors that aligns with the intrinsic shape of the structure. The first principal component (PC1) describes the direction of the largest variance and the length of the structure. The second principal component (PC2) describes the direction of the largest variance orthogonal to PC1 and the width of the structure. The third principal component (PC3) describes the direction of the largest variance orthogonal to PC1 and PC2, and the height of the structure.
We determined the structure axis lengths by projecting the data onto the principal components and measuring the extent between the minimum and maximum, thereby providing an upper bound. For the length, this is a good approximation, but due to the irregular shape of the clouds, defining the height and width leads to a large overestimation in the majority of the cases.
@@ 3.2.3 Orientation calculation
The structure orientation is determined by the angle between PC1 and a chosen reference axis. This angle is calculated by applying the arc cosine to the absolute value of the cosine similarity between the two vectors, and then converting the result to degrees. For consistency and easier comparison, all structures are defined as oriented in the positive direction relative to the reference axis, with angles measured clockwise from it. The angle between PC1 and the Y-axis (face-on) is defined as $\alpha$, while $\beta$ represents the angle between PC1 and the Z-axis (edge-on). Circular statistics are applied to calculate the mean and standard deviation of all angles $\alpha$ and $\beta$. The angles are weighted by mass to consider each structure’s relative importance.
@@ 3.2.4 Undulation fitting
where $z$ corresponds to the projected Y-coordinate, obtained by rotating the superclouds counterclockwise by their average angle $\alpha$ to lie in the plane, and $f(z)$ gives the corresponding Z-position. The parameter $a$ is the amplitude, $\omega$ the angular frequency, $\phi$ the phase shift, and $d$ the vertical offset. The function includes an exponential modulation of both amplitude and frequency: $\epsilon_{a}$ describes the exponential growth or decay of the amplitude along the wave, while $\epsilon_{\omega}$ describes a linear change in frequency. Positive values for $\epsilon_{a}$ indicate a growth in amplitude over distance, while negative values indicate damping. For $\epsilon_{\omega}$, a positive value indicates a growth in frequency, resulting in a decrease in wavelength via the relation $\lambda=2\pi/\omega$, where $\lambda$ is the wavelength and $\omega$ the angular frequency.
The fitting approach is simplified to give an initial characterization of the underlying undulation and should not be interpreted as implying a physical oscillation.
@@ 4.1 HOP cloud catalog
HOP identifies 40 clouds in the Edenhofer2024 dust map, visualized in the left panel of Fig. 1. The cloud masses range from $3.8\times 10^{4}$ to $2.8\times 10^{6}$ M, a selection of the calculated cloud properties including the mass and orientation can be found in Table 2. All clouds with their corresponding ID numbers are visualized in Fig. 7.
Figure 2: Orientation distribution of the 40 HOP clouds. Colors indicate the same clouds as in Fig. 1, the distance from the center reflects the cloud mass, and the solid line represents the mean orientation. Top: Face-on view from the North Galactic Pole, with an average angle of $\alpha=22.6\pm 25.5\degr$. Bottom: Edge-on view from the Galactic Center, with an average angle of $\beta=92.3\pm 23.4\degr$.
@@
Figure 3: Face-on view of the HOP clouds color-coded by their mean altitude in Z. The Sun is represented by the yellow dot in the center, and the Galactic Center is located to the right.
@@ 4.2 Supercloud catalog
Taking the orientation, proximity, and 3D alignment of the HOP clouds into account, we identify seven superclouds, visualized in the left panel of Fig. 4. The corresponding HOP cloud components are listed in Table 1, with the supercloud ID numbers visualized in Fig. 8.
To connect the superclouds, the first criterion is the face-on orientation of the HOP clouds, defined by $\alpha$ and shown in the right panel of Fig. 1. The primary motivation for grouping these structures lies in their coherent orientation. If adjacent clouds show a similar orientation, defined as differing by no more than 35°, they are considered as initial candidates for a supercloud. This threshold is chosen to connect coherent, large-scale structures while avoiding random overlaps.
We then verified the preliminary connections by taking the spatial distribution of the clouds into account. This ensures that the clouds grouped by orientation form a consistent, physical structure. Candidate clouds that share similar orientations, are spatially connected, and appear as a single extended feature are confirmed as parts of the superclouds. Clouds that deviate from the group orientation, yet still show a spatial connection, are also included as candidate clouds. This case of spatial connectivity is only applied when neighboring clouds clearly form a continuous feature.
It is likely that external forces have played a role in the fragmentation of the superclouds during their lifetime. Clouds showing no connection to the supercloud candidates are verified concerning their general 3D alignment. If displaced clouds follow the spatial trend of a larger, already verified group, they are accepted candidates based on their 3D alignment. If these clouds cannot be assigned to another candidate group following the mentioned steps, they are confirmed as part of the first suggested supercloud.
The connection of the superclouds is re-verified by comparing the final segmentation to the parameter test runs with different lower mincells parameters, ensuring that potential links via smaller clouds are not missed.
Figure 4: Face-on view of the seven superclouds defined in the local Milky Way (left) and their corresponding orientation (right). Each cloud is assigned a distinct color. The length of the line corresponds to the calculated length of the cloud. The Sun is represented by the yellow dot in the center, and the Galactic Center is located to the right. An interactive 3D version of the seven superclouds is available here.
@@ 4.2.1 Supercloud identification
For the more disrupted orange and yellow cloud group, we adopt the name Malpolon Cloud. Its irregular morphology and puffed-up appearance in Z is likely caused by the influence of nearby high-mass stars, as traced by zari2023 and the ALS III Catalog of Galactic OB stars (pantaleoni2025), which is a recent expansion and Gaia DR3 update of the catalog in ALSII.
The green supercloud, which we name the Vela Ridge Cloud, is the least spatially coherent among the identified superclouds, with its components fragmented and widely separated. A possible explanation is that the Vela Ridge Cloud itself is not an independent supercloud but rather a disrupted part of the SSE, as orbital tracebacks of young star clusters hint that past clustered supernovae may have occurred in this region (Swiggum et al., in prep). This theory of a past connection is also supported by the presence of high-mass stars in the canyon that separates clouds 4 and 33 from cloud 2. The stellar component extends the Spur further to almost double the initially defined length, linking the separated cloud components (pantaleoni2025). Spatially, it is not unlikely that clouds 29 and 39 (referred to as the L-clouds) are being sheared off from the SSE due to external forces, such as feedback from high-mass stars in this region (zari2023; pantaleoni2025), or supernovae from the Messier 6 cluster family that partially occupies this region (swiggum_most_2024). Additionally, clouds 23 and 26 may have originally been part of the Split, which is supported by their physical connection. The displacement is likely a result of the event responsible for the ring they surround. We kept the Vela Ridge Cloud as a supercloud in this work, as we cannot verify any of the alternative interpretations yet.
Lastly, the brown supercloud, located at the upper edge of the map, is given the name Anguis Cloud. While the two supercloud components are not connected, comparing our results with the extended version of the Edenhofer2024 dust map and the dust map by vergely2022 shows that they are likely part of a supercloud extending beyond the map.
@@ 4.2.2 Supercloud properties
All superclouds we identify are truncated by the artificial boundary of the 3D reconstruction and are therefore incomplete. Their cloud components are either cut off at the border of the map or show potential extensions when compared to the extended 2 kpc Edenhofer2024 dust map and the dust map from vergely2022. These maps are not used for the segmentation itself, due to their decreasing reliability beyond the boundary of the main map and their lower resolution. Instead, they serve as additional references for assumptions on the possible continuation of the superclouds.
As shown by alves2020, the Radcliffe Wave extends to the Cygnus X complex. zucker2020 calculated accurate distances for SFRs in the Star Formation Handbook (reipurth2008_sfr_north; reipurth2008_sfr_south) up to 2.5 kpc from the Sun. Comparing their results with the superclouds, we find that the Natrix Cloud is likely to extend to the Rosette Nebula, while the SSE connects to M20, Lagoon, L291, NGC6604, M17, M16, and L379. It is interesting to note that each supercloud, excluding the Anguis Cloud, contains at least a few known SFRs. This absence could indicate a less active or earlier phase of cloud evolution, but it might also result from observational restraints or incomplete coverage of the region.
The masses of the inferred superclouds range from $1.1$ to $3.6\times 10^{6}$ M, with the Anguis Cloud as an outlier at $7.8\times 10^{5}$ M. All superclouds, except the Anguis Cloud, show lengths ranging from approximately 1.5 to 2.5 kpc, which is calculated using PCA and taking all cloud components into account. A selection of the calculated cloud properties, including mass and orientation, can be found in Table 3. Their mean orientation is characterized by a face-on pitch angle $\alpha=33.5\pm 4.0\degr$, and edge-on angle $\beta=87.7\pm 4.2\degr$, making an association with a major spiral arm unlikely (Franco2002; efremov2009). We show the face-on orientation of the superclouds in the right panel of Fig. 4. The identified superclouds are consistent with the characteristic scales and masses predicted in earlier studies (Elmegreen1987; park2023), representing the largest scale of ISM structures in the local Milky Way and likely throughout the Galaxy as a whole.
In total, the Edenhofer2024 dust map has a gas mass of $3.3\times 10^{7}$ M. For Z = ± $\pm\,100$ pc, the volume filling factor of the superclouds is 9.94%, containing 55.95% of the gas mass in the Galactic plane.
@@ 4.2.3 Undulation pattern
Notably, the undulation observed in the Radcliffe Wave is not unique. All but one of the seven superclouds show an undulating pattern in the Z/Y plane, indicated by the color pattern of Fig. 3, and are likely oscillating (konietzka2024). Here, the upper part of the Radcliffe Wave appears flat in Z, as a result of averaging the change in height across the entire length of HOP cloud 0, erasing the underlying pattern.
In Fig. 5 we show the sinusoidal fit for the selected superclouds: the Malpolon Cloud, Natrix Cloud, Radcliffe Wave, and Vela Ridge Cloud. The SSE and the Anguis Cloud are excluded from the analysis, as they show a vertical displacement, but it is not as prominent as for the previous examples. The Split is also excluded, as it appears unusually flat, with only two noticeable downward extensions at Y = -200 pc (referred to as the ”C” and discussed in edenhofer2024_c) and Y = 500 pc. The three superclouds excluded from the fitting can be seen in Fig. 9. All resulting fit parameters are listed in Table 4, the fitting approach is simplified to give an initial characterization of the underlying undulation.
For the Malpolon Cloud, we excluded the three highest altitude clouds (clouds 10, 16, and 17) while fitting. It is likely that these clouds were pushed up by nearby high-mass stars (zari2023; pantaleoni2025), and therefore distort the inherent structure of the supercloud. The Malpolon Cloud, Natrix Cloud, and Radcliffe Wave are well described by a damped sinusoidal wave. For the Malpolon Cloud, the parameters indicate decay in the amplitude and growth in the wavelength. Both damping parameters for the Natrix Cloud indicate growth over distance. For the Radcliffe Wave, the parameters show decay in amplitude and wavelength.
The Vela Ridge Cloud is fitted with a simple sinusoidal function. While the undulation is visible, the large wavelength is likely a consequence of the fragmented nature of the supercloud. This, together with the incompleteness, supports the idea that Vela Ridge is not an individual supercloud.
Figure 5: 2D projection in the Z/Y plane of the selected superclouds with a sinusoidal fit. For the fit, the clouds are rotated counterclockwise by their face-on angle ($\alpha$) in the X/Y plane. The first three clouds are fit with a damped sinusoidal, the fourth cloud with a simple sinus function. All fit parameters are listed in Table 4.
@@ 5 Discussion
A key implication of the main results in this paper is that these superclouds are unlikely to be associated with the spiral arms of the Milky Way, which have significantly smaller pitch angles (on average 10, as shown in reid2019). This includes the most massive of the local superclouds, the Radcliffe Wave, which is normally associated with the Local Arm (also known as the Orion Arm or the Orion Spur; see, e.g., swiggum2022). Instead, the local superclouds seem to be part of a distinct, coherent structure that may reflect larger-scale dynamical processes in the Milky Way, perhaps akin to the high-pitch-angle gas lanes observed in other galaxies.
@@
Additionally, given that none of the seven superclouds seem to be associated with a Milky Way spiral arm, this implies that there is no obvious external trigger for the formation of GMCs inside of them, like a spiral shock. An initial comparison with the spatial distribution of known star-forming GMCs shows that nearly all lie within these superclouds, with many concentrated along their central regions of higher density, a result aligning with earlier results from Dame1986 and Elmegreen1987. This supports a molecular cloud formation scenario in the local Milky Way in which H2 forms in the denser regions of the neutral atomic cold medium without an obvious external trigger mechanism. However, this interpretation is tentative, and more work is needed to qualify this statement. Since GMCs are the primary sites of star formation, the superclouds presented in this work are likely to play an indirect but critical role in shaping the star formation process in the local Milky Way.
@@
Of the seven superclouds identified in this study, only two were previously known: the Split (Lallement2019) and the Radcliffe Wave (alves2020). While the Split lies on the Galactic midplane, the Radcliffe Wave surprisingly undulates above and below it. Recent work by konietzka2024, using young stellar clusters, confirmed that the undulations of the Radcliffe Wave represent vertical oscillations relative to the plane. Our results show that such undulations are not unique to the Radcliffe Wave. In fact, all superclouds, except the Split, exhibit similar vertical displacements. The origin of these undulations remains uncertain. One possibility is that they resulted from shocks or gravitational perturbations as superclouds interacted with spiral arms in the past, but further investigation is needed to test this scenario.
@@
The identified superclouds can be associated with multiple known structures in the solar neighborhood. Spatially, the Anguis and the Malpolon Cloud appear to be a small section of the Cepheus Spur (ALSII). This spur extends from Cygnus OB2 toward the Galactic anticenter, as traced by young stellar populations and in more extended dust maps (e.g., Lallement2019). The bottom part of the Radcliffe Wave (clouds 27, 34, 30, 12, and 18) and of the Vela Ridge Cloud (clouds 21 and 2) define the boundaries of the GSH 238+00+09 supershell (heiles1998). This kiloparsec-long dust shell likely formed from clustered supernovae associated with the Cr135 cluster family (swiggum_most_2024), and occupies the full width between the two superclouds. Recently, soler2025 calculated the kinetic energy densities of the local ISM based on the Edenhofer2024 dust map. Kinetic energy overdensities represent areas that differ roughly by a factor of ten or more from the mean value. The largest of these overdensities can be connected to the Radcliffe Wave and the Split, likely as a result of past supernovae or large-scale galactic dynamics.
@@
Figure 6: Properties calculated for the seven superclouds. The colors have the same meaning as in Fig. 4. RadWave refers to the Radcliffe Wave. Top: Total mass of the superclouds. Second: Total length of the superclouds. Third: Linear mass of the superclouds, the ratio of total mass to total length. Bottom: Average density of the superclouds. The trends represent a lower bound, as all superclouds are incomplete and likely continue beyond the map.
The well-defined, quasi-parallel pattern of superclouds shown in Fig. 4 suggests a relatively undisturbed local ISM. While stellar feedback is clearly reshaping the gas on kiloparsec scales, for example, in the case of the Heiles supershell (swiggum_most_2024), it does not appear to have significantly disrupted the overall supercloud pattern. Based on the formation scenario for GMCs as discussed above, we can make a conservative, lower age estimation by considering the ages of the oldest stars formed within the superclouds. For the Radcliffe Wave, this corresponds to approximately 20-30 Myr. Compared to kinematic timescales, this lower age estimate is short. Recent measurements of the local ISM velocity dispersion, approximately $11\,\mathrm{km\,s^{-1}}$ (see soler2025, and references therein), suggest that a perturbation would require over $90\,\mathrm{Myr}$ to cross the full length of a supercloud. This timescale is about three times longer than the estimate based on stellar ages and exceeds the typical lifetimes of CO clouds (see, e.g., mivilledeschenes2017; chevance2020). These discrepancies imply that superclouds are unlikely to form by a bottom-up process, i.e., from small to large scales, as would be expected in a scenario driven purely by stellar feedback. Instead, large-scale Galactic dynamics and instabilities probably play a central role in the formation of the kiloparsec-scale density structures we identify as superclouds (see, e.g., kim2002; kim_ostriker_tigress2017; smith2020; colman2022). This does not exclude the influence of energy and momentum injection from interior star formation (swiggum_most_2024), but it points to the need for additional physical processes to explain the observed coherence and scale of these structures.